# HARec: Hyperbolic Graph-LLM Alignment for Exploration and Exploitation in Recommender Systems

### v1

@@ 4.3. Hyperbolic Alignment
(15) $s_{i}$ $=\exp_{x}(e_{i})$
$s^{\prime}_{i}$ $=\textit{MLP}(s_{i})$
where $\exp_{x}$ is defined as in 3, $e_{i}$ represents the semantic embeddings from the encoder, $s_{i}$ is the dimension-aligned semantic embedding, and $s^{\prime}_{i}$ denotes the final projected semantic embedding in hyperbolic space. The hyperbolic alignment loss is then calculated as:
(16) $\ell_{align}(i)=d_{\mathcal{H}}^{2}(h_{i},\textbf{s}^{\prime}_{i})$
@@
4.4. Gradient Analysis of Semantic Alignment.
To further explore the impact of aligning semantic embeddings within the hyperbolic model, we analyze the gradient behavior with respect to nodes of varying norms in Proposition 4.1. This provides insight into the hierarchical preservation properties of the hyperbolic space, which are absent in Euclidean space.
@@
The gradient magnitude of semantic alignment in hyperbolic space $\|\nabla_{\mathbf{x}}d_{\mathcal{H}}\|\approx\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}$ enables adaptive updates for hierarchical preservation, while Euclidean space maintains constant magnitude $\|\nabla_{\mathbf{x}}d_{E}\|=1$, where the $d_{\mathcal{H}}$ and $d_{\mathcal{E}}$ represent the hyperbolic distance and Euclidean distance, respectively.
Hyperbolic Space. Suppose $\mathbf{x}$ is a hyperbolic embedding derived from  4.1, and $\mathbf{y}$ is the semantic embedding from  4.2. The hyperbolic distance is given by:
(17) $d_{\mathcal{H}}(\mathbf{x},\mathbf{y})=\operatorname{arccosh}(z)$
where $z=-\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}$. For large $\|\mathbf{x}\|$ and $\|\mathbf{y}\|$, given $x_{0}=\sqrt{1+\|\mathbf{x}\|^{2}}$ and $y_{0}=\sqrt{1+\|\mathbf{y}\|^{2}}$, we adopt the following approximations for further calculation:
(18) $x_{0}\approx\|\mathbf{x}\|,\quad y_{0}\approx\|\mathbf{y}\|,\quad\sqrt{z^{2}-1}\approx z$
(19) $z=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}\approx\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)$
Let $x_{j}=\|\mathbf{x}\|\hat{x}_{j}$ and $y_{j}=\|\mathbf{y}\|\hat{y}_{j}$ (where $\hat{x}_{j}$ represents the direction of $x_{j}$). The partial derivative of $z$ with respect to $x_{j}$ is:
(20) $\frac{\partial z}{\partial x_{j}}=\frac{x_{j}}{x_{0}}y_{0}-y_{j}\approx\frac{x_{j}}{\|\mathbf{x}\|}\|\mathbf{y}\|-y_{j}=\|\mathbf{y}\|(\hat{x}_{j}-\hat{y}_{j})$
Thus, the gradient norm of $d_{\mathcal{H}}$ for $\mathbf{x}$ can be approximated by:
(21) $\|\nabla_{\mathbf{x}}d_{\mathcal{H}}\|=\frac{\|\nabla_{\mathbf{x}}z\|}{\sqrt{z^{2}-1}}\approx\frac{\|\nabla_{\mathbf{x}}z\|}{z}=\frac{\|\mathbf{y}\|\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)}=\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}$
This result indicates that nodes with large norms undergo smaller gradient updates, preserving local structures, while nodes with smaller norms adjust more, refining global hierarchical relationships and capturing hierarchical structure effectively.
Euclidean Space. For comparison, we examine the gradient behavior in Euclidean space. The Euclidean distance is defined as:
(22) $f(\mathbf{x})=d_{E}(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|$
The gradient of $f(\mathbf{x})$ with respect to $\mathbf{x}$ is:
(23) $\nabla_{\mathbf{x}}f(\mathbf{x})=\frac{\mathbf{x}-\mathbf{y}}{\|\mathbf{x}-\mathbf{y}\|}=\frac{\mathbf{x}-\mathbf{y}}{d_{E}(\mathbf{x},\mathbf{y})}$
which has a magnitude of:
(24) $\|\nabla_{\mathbf{x}}f(\mathbf{x})\|=1$
In summary, the analysis shows that hyperbolic space provides adaptive gradient magnitudes that depend on the norms of the nodes, effectively capturing and preserving hierarchical structures. In contrast, Euclidean space, with its constant gradient magnitude, cannot adapt to hierarchical nuances, making it less suitable for modeling hierarchical data in recommender systems.
4.5. Hyperbolic Hierarchical Structure
Figure 3. Hierarchical Structure Visualization.
In this subsection, we introduce a mechanism that allows users to control the level of exploration in their recommendations by constructing a hierarchy tree based on the learned hyperbolic embeddings, an example is shown in Figure 3. These embeddings integrate both collaborative information and language semantics, enabling more nuanced recommendations.
@@ 4.5. Hyperbolic Hierarchical Structure
Hierarchy Tree. The primary motivation for constructing a hierarchy tree is to uncover hidden structures in the data that are not explicitly reflected in the dataset. For instance, consider the Amazon-CD dataset: a CD might be categorized under genres like “folk music”, but it cannot be labeled as “folk music preferred by teenagers” or further refined to “folk music preferred by teenagers who live in large cities”. However, in reality, such nuanced preferences exist within the collaborative information. When recommending based solely on past interactions, we may overlook these hidden layers of structure that are implicit in the dataset (e.g., “music” → “folk music” → “folk music preferred by teenagers” → “folk music preferred by teenagers who live in large cities”).
As previously discussed, in hyperbolic space, the norm of the embeddings reflects the level of exploration, where depth search facilitates exploitation and breadth search enables exploration. Building on this concept, we use the learned embeddings to form the bottom layer of the hierarchy tree. We then apply K-Means clustering in hyperbolic space to derive higher-level cluster nodes, which act as pseudo clustering nodes within the hierarchy, representing group preferences. For example, if an item is classified as “folk music preferred by teenagers who live in large cities”, its parent cluster may represent “folk music preferred by teenagers”, while its sibling clusters could include “folk music preferred by teenagers who live in small towns”. This hierarchical clustering process continues iteratively until reaching the root, thereby creating a multi-level tree structure that captures different levels of exploration.
@@
Exploration-Exploitation Balance. To allow users control over the balance of exploration and exploitation, we introduce two adjustable parameters: temperature $\tau$ and hierarchy level $l$. In the standard recommendation approach, items are recommended based on similarity to the user’s embedding, which may limit exploration. For users with focused preferences, this method may consistently recommend items that align too closely with their past choices, offering little opportunity to explore different categories.
@@
To address this issue, we modify the recommendation process by replacing a portion of the recommended items with alternatives from other branches of the hierarchy tree. This modification gives users control over the balance between exploration and exploitation. Temperature $\tau$ allows users to specify the proportion of original recommendations to replace, while level $l$ determines the extent of exploration within the hierarchy. For example, if a user selects $\tau=0.5$ and $l=2$, half of the originally recommended items will be retained, while the other half will be replaced by items sampled from nodes two levels up in the hierarchy tree, with selections made from the connected bottom-layer nodes.
5. EXPERIMENTS
@@
Table 1. Comprehensive model comparison based on utility and diversity metrics, where higher values signify superior performance. The best-performing results for each metric are highlighted in bold, with the second-best results underlined.
@@
Table 2. Comparison with feature-enhanced Euclidean baselines, with the best-performing results highlighted in bold.
@@
5.1. Experimental Settings
Datasets. To evaluate our proposed model, we utilize three widely-used public datasets: Amazon-books, which contains user purchase behaviors within the book category on Amazon; Yelp, which records customer ratings and reviews for restaurants on Yelp; and Google-reviews, which includes user reviews and business metadata from Google Maps. Detailed statistics for these datasets , including the distribution of head and tail items, are provided in Table 3. The high proportion of tail users (T80) across these datasets highlights the prevalence of long-tail distributions, supporting the assumption of a power-law distribution in real-world data.
Evaluation Metrics. We assess our recommended items using both utility and diversity metrics, demonstrating that our model achieves precise recommendations for users’ preference while enhancing diversity in the recommendation list.
For utility, we adopt two widely-used ranking metrics: Recall and NDCG, which measure model effectiveness. To evaluate diversity, we employ three metrics: Distance Diversity (Ziegler et al., 2005; Zhang and Hurley, 2008), Shannon Entropy (Mehta et al., 2012), and Expected Popularity Complement (Vargas and Castells, 2011).
(25) $Div(u)=\frac{\sum_{i\in R_{u}}\sum_{j\in R_{u},j\neq i}d(i,j)}{\text{Number of}(i,j)\text{ pairs}}$
(26) $H=-\sum_{i\in\text{set}(\sum_{u}R_{u})}p(i)\log p(i)$
(27) $EPC=1-\frac{1}{|\text{set}(\sum R_{u})|}\sum_{i\in\text{set}(\sum_{u}R_{u})}\frac{\text{pop}(i)}{\max_{i}(\text{pop}(i))}$
Baselines. We compare our model with state-of-the-art baselines in Euclidean and Hyperbolic spaces. Euclidean models include:
  • LightGCN (He et al., 2020): A lightweight graph recommender model by removing redundant neural modules in graph convolution.
  • TAG-CF (Ju et al., 2024): Enhances collaborative filtering by performing a single message-passing step only at inference time.
  • LightGCL (Cai et al., 2023): Uses singular value decomposition (SVD) for robust graph augmentation without data augmentations.
  • SimGCL (Yu et al., 2022): Replaces graph augmentations in contrastive learning by introducing random noise during initialization.
  • HCCF (Xia et al., 2022): Integrates hypergraph and contrastive learning, capturing both local and global user dependencies.
Hyperbolic GNN models include:
  • HGCF (Sun et al., 2021): Combines hypergraph learning with collaborative filtering using hyperbolic margin ranking loss.
  • HICF (Yang et al., 2022a): Enhances head and tail item performance with hyperbolic geometry, improving utility for long-tail items.
Implementation Details. We set embedding dimensions to 50 and a training batch size of 1024. Embeddings in hyperbolic space are optimized using Riemannian SGD (RSGD) (Bécigneul and Ganea, 2018; Sun et al., 2021), while the MLP adapter is optimized with Adam (Kingma, 2014). Training incorporates early stopping technique based on Recall@10 with 10 patience steps. For the hierarchical structure, we use a $1:8$ ratio for parent and child nodes. We generate user and item profiles with GPT-3.5-turbo and encode them with text-embedding-ada-002 (Neelakantan et al., 2022) for Amazon-books and Yelp, and BERT (Devlin, 2018) for Google-reviews. The profile decoder backbone is Llama 3.1-8B. All experiments are conducted on NVIDIA A40/A100 GPUs with 40GB memory.
Table 3. Statistics of the experimental data.
@@ 5.1. Experimental Settings
5.2. Performance Comparison
Overall Comparison. The performance comparison across all models is presented in Table 1. Our HARec consistently outperforms the baselines in both utility and diversity metrics, achieving up to $5.49\%$ improvement in utility metrics and $11.39\%$ increase in diversity metrics. Notably, most baselines struggle to achieve strong performance in both utility and diversity simultaneously. For instance, while HICF performs second-best in utility metrics on the Amazon-books dataset, outperforming other baselines by a significant margin (excluding our model), it lacks the same advantage in diversity metrics, where several baselines achieve better results. This highlights the superiority of our model, which achieves state-of-the-art performance in both utility and diversity: a feat not previously attained by any baseline.
Baselines with Feature Enhancement. To ensure a fair comparison, we further enhance Euclidean baselines by integrating feature information. Specifically, we align semantic embeddings with the original Euclidean embeddings following the structure in Section 4.2, but in Euclidean space. The results are presented in Table 2, with enhanced baselines denoted by “+”. The data shows that our HARec still surpasses all baselines in both utility and diversity. Interestingly, while the feature-enhanced versions of the Euclidean baselines generally improve in utility metrics, they do not exhibit a stable improvement in diversity metrics. This aligns with our theoretical analysis in Section 4.5, which demonstrates the advantages of our alignment model in the hyperbolic space.
Table 4. Layer-wise norms within the hierarchy structure.
Data level-0 level-1 level-2 level-3 level-4
Amazon-books 0.2388 0.4859 0.5143 0.5889 0.6773
Yelp 0.1758 0.3500 0.4719 0.5590 0.6307
Google-reviews 0.4374 0.8585 1.0682 1.1731 1.2479
Figure 4. Ablation study on model variants.
Table 5. Performance on H20 (head) and T80 (tail) items, with the top hyperbolic results highlighted in bold. $\Delta_{\mathcal{H}}(\%)$ represents the improvement over the second-best hyperbolic baseline, which is underlined.
Data Amazon-books Yelp Google-reviews
@@ 5.2. Performance Comparison
Figure 5. Performance across different levels of exploration.
5.3. Ablation Study

### v2

@@ 4.3 Hyperbolic Alignment
$s_{i}=\exp_{x}(e_{i}),\quad s^{\prime}_{i}=\textit{MLP}(s_{i})$ (13)
where $\exp_{x}$ is defined as in Equation 3, $e_{i}$ represents the semantic embeddings from the encoder, $s_{i}$ is the dimension-aligned semantic embedding, and $s^{\prime}_{i}$ denotes the final projected semantic embedding in hyperbolic space. The hyperbolic alignment loss is then calculated as:
$\ell_{align}(i)=d_{\mathcal{H}}^{2}(h_{i},s^{\prime}_{i})$ (14)
@@
4.4 Gradient Analysis of Hyperbolic Alignment
To further explore the impact of aligning semantic embeddings within the hyperbolic model, we analyze the gradient behavior with respect to nodes of varying norms in Proposition 4.1, proof can be found in Appendix A. This provides insight into the hierarchical preservation properties of the hyperbolic space, which are absent in Euclidean space.
@@ 4.4 Gradient Analysis of Hyperbolic Alignment
Let $\mathbf{x},\mathbf{y}$ denote embeddings in hyperbolic space $\mathcal{H}^{n}$ with corresponding unit direction vectors in space-like dimension $\hat{\mathbf{x}},\hat{\mathbf{y}}$, and let $\theta$ be the angle between them. The gradient magnitude of semantic alignment satisfies:
$\|\nabla_{\mathbf{x}}d_{\mathcal{H}}(\mathbf{x},\mathbf{y})\|\approx\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}.$ (15)
@@
This facilitates adaptive gradient updates that intrinsically preserve hierarchical structures. In contrast, Euclidean space exhibits constant gradient magnitude $\|\nabla_{\mathbf{x}}d_{\mathcal{E}}(\mathbf{x},\mathbf{y})\|=1$, where $d_{\mathcal{H}}$ and $d_{\mathcal{E}}$ denote hyperbolic and Euclidean distances respectively.
This proposition highlights that hyperbolic space offers adaptive gradient magnitudes based on node norms. Nodes with large norms (fine-grained preferences, small positional changes cause significant distance changes) receive smaller gradient updates, ensuring precise distance adjustments and preserving local structures. Meanwhile, nodes with smaller norms (abstract preferences, positional changes have less impact on distances) adjust more, refining global relationships and effectively capturing the hierarchical structure.
4.5 Hierarchical Representation Structure
In this subsection, we design a mechanism that allows users to balance the exploration-exploitation trade-off by constructing a hierarchical representation structure based on the learned hyperbolic embeddings.
@@ 4.5 Hierarchical Representation Structure
Hierarchy Tree. The primary motivation for constructing a hierarchy tree is to uncover hidden structures in the data that are not explicitly reflected in the dataset. For instance, consider the Amazon-CD dataset: a CD might be categorized under genres like “folk music”, but it cannot be labeled as “folk music preferred by teenagers”. However, in reality, such nuanced preferences exist within the collaborative information. When recommending based solely on past interactions, we may overlook these hidden layers of structure that are implicit in the dataset (e.g., “music” → “folk music” → “folk music preferred by teenagers”).
To address this, we introduce a hyperbolic hierarchical clustering method based on hyperbolic embeddings to build the hierarchy tree. Specifically, we iteratively group embeddings into clusters and use the centroid of each cluster as the data point to be clustered in the upper layer, referred to as a pseudo-cluster node. However, the number of cluster nodes in each layer is not predefined and must be determined. Inspired by Dasgupta’s cost (Dasgupta, 2016), which encourages pushing edge cuts as far down the tree as possible, the optimal tree is required to be binary. Therefore, we fix the proportion of cluster nodes between layers as $k=2$. The details of the algorithm are presented in Algorithm 1.
Algorithm 1 Hyperbolic Hierarchical Clustering
  Input: Hyperbolic embeddings $X$, layer proportion $k$
  Set the maximum number of layers $L=\log_{k}(|X|)$
  Initialize the points of bottom layer $D_{L}=X$
  for $l=L$ to $1$ (bottom to top) do
     Randomly initialize cluster centroids in the hyperbolic space $M_{l}=\{\mu_{l}\ |\ \mu_{l}\in\mathbb{H}\}$, $|M_{l}|=|D_{l}|/k$
     repeat
        for each point $d_{l,i}\in D_{l}$ do
           Assign $d_{i}$ to the cluster with the nearest centroid:
           $c_{l,i}=\arg\min_{\mu\in M_{l}}d_{\mathcal{H}}(d_{i},\mu)$
        end for
        for each cluster $c_{l,i}\in C_{l}$ do
           Update the centroid $\mu_{l,i}$ as the hyperbolic mean $m_{\mathcal{H}}$ of all points assigned to cluster $c_{l,i}$:
           $\mu_{l,i}=m_{\mathcal{H}}(\{d\ |\ d\in C_{l,i}\})$
        end for
     until Centroids $M_{l}$ do not change (convergence)
     Initialize the points of the upper layer $D_{l-1}=M_{l}$
  end for
@@
Exploration-Exploitation Balance. To allow users to balance the exploration-exploitation trade-off, we introduce two adjustable parameters: temperature $\tau$ and hierarchy layer $l$. In the standard recommendation approach, items are recommended based on similarity to the user’s embedding, which may limit exploration. For users with focused preferences, this method may consistently recommend items that align too closely with their past choices, offering little opportunity to explore different categories.
@@
To address this issue, we modify the recommendation process by replacing a portion of the recommended items with alternatives from other branches in the hierarchy tree. This modification gives users control over the balance between exploration and exploitation. Temperature $\tau$ allows users to specify the proportion of original recommendations to replace, while layer $l$ determines the starting point for exploration. For example, suppose a user selects $\tau=0.5$ and $l=5$. In that case, half of the original recommendations are retained, and the other half is replaced with items sampled from its ancestor cluster at layer $5$ of the hierarchy tree.
Complexity Analysis. To demonstrate the scalability and efficiency of our approach, we analyze the complexity of constructing and applying the hierarchical representation structure. Let $N$ denote the total number of users and items, and let $k$ be the number of recommendations per user. The hierarchy tree is built using iterative $k$-means clustering in hyperbolic space, where cluster nodes are assigned only to real node embeddings for efficiency.
First, computing and storing pairwise distances requires $O(N^{2}/2)$ operations. Each $k$-means iteration has complexity $O(i)$, and with $O(\log N)$ hierarchical levels, the total complexity of constructing the hierarchy tree is:$O\left(\frac{N^{2}}{2}+i\log N\right)$. For recommendation retrieval, finding the top-$k$ recommendations involves nearest neighbor search in hyperbolic space, which runs in $O(k\log N)$. Adjusting recommendations by retrieving ancestor clusters at layer $l$ requires only tree traversal, which runs in $O(k)$.
Overall, our hierarchical framework ensures efficient scaling while maintaining flexibility in recommendation adjustments. This design enables HERec to dynamically balance exploration and exploitation while ensuring scalability for large-scale recommender systems.
5 EXPERIMENTS
@@
Table 1: Comprehensive model comparison based on utility and diversity metrics, where higher values signify superior performance. The best-performing results for each metric are highlighted in bold, with the second-best results underlined.
@@
Table 2: Comparison with feature-enhanced Euclidean baselines, with the best-performing results highlighted in bold.
@@
Table 3: Statistics of the experimental data.
@@
Table 4: Performance on H20 (head) and T80 (tail) items, with the top hyperbolic results highlighted in bold. $\Delta_{\mathcal{H}}(\%)$ represents the improvement over the second-best hyperbolic baseline, which is underlined.
Datasets Amazon-books Yelp Google-reviews
@@
5.1 Experimental Settings
Datasets. To evaluate our proposed model, we utilize three widely-used public datasets: Amazon-books, which contains user purchase behaviors within the book category on Amazon; Yelp, which records customer ratings and reviews for restaurants on Yelp; and Google-reviews, which includes user reviews and business metadata from Google Maps. Detailed statistics, including the distribution of head and tail items, are provided in Table 3. The high proportion of tail users (T80) highlights the prevalence of long-tail distributions, supporting the assumption of a power-law distribution in real-world data.
Evaluation Metrics. We assess our recommended items using both utility and diversity metrics, demonstrating that our model achieves precise recommendations for users’ preference while enhancing diversity in the recommendation. For utility, we adopt two widely used ranking metrics: Recall and NDCG, which measure the effectiveness of the model. To evaluate diversity, we employ three metrics: Distance Diversity (Div) (Ziegler et al., 2005; Zhang & Hurley, 2008), Shannon Entropy (H) (Mehta et al., 2012), and Expected Popularity Complement (EPC) (Vargas & Castells, 2011).
$Div(u)=\frac{\sum_{i\in R_{u}}\sum_{j\in R_{u},j\neq i}d(i,j)}{\text{Number of}(i,j)\text{ pairs}}$ (16)
$H=-\sum_{i\in\text{set}(\sum_{u}R_{u})}p(i)\log p(i)$ (17)
$EPC=1-\frac{1}{|\text{set}(\sum R_{u})|}\sum_{i\in\text{set}(\sum_{u}R_{u})}\frac{\text{pop}(i)}{\max_{i}(\text{pop}(i))}$ (18)
Baselines. We compare our model with state-of-the-art baselines in both Euclidean and Hyperbolic spaces:
  • LightGCN (He et al., 2020): A lightweight graph convolution neural network model for recommendation by removing redundant neural modules in graph convolution.
  • TAG-CF (Ju et al., 2024): A test-time augmentation framework that enhances recommendation by performing a single message-passing step only at inference time.
  • LightGCL (Cai et al., 2023): A contrastive learning model that uses singular value decomposition (SVD) for robust graph augmentation without data augmentation.
  • SimGCL (Yu et al., 2022): A contrastive learning model that replaces graph augmentations in contrastive learning by introducing random noise during initialization.
  • HCCF (Xia et al., 2022): A collaborative filtering framework that integrates hypergraph and contrastive learning, capturing both local and global user dependencies.
  • HGCF (Sun et al., 2021): A hyperbolic GNN model that combines hypergraph learning with collaborative filtering using hyperbolic margin ranking loss.
  • HICF (Yang et al., 2022a): A hyperbolic recommendation model that enhances head and tail item performance with hyperbolic geometry, improving utility for long-tail items.
Training Details. We set embedding dimensions to 50 and a training batch size of 1024. Embeddings in hyperbolic space are optimized using Riemannian SGD (RSGD) (Bécigneul & Ganea, 2018; Sun et al., 2021), while the MLP adapter is optimized with Adam (Kingma, 2014). Training incorporates early stopping technique based on Recall@10 with 10 patience steps. We generate user and item profiles using GPT-3.5-turbo and encode them with text-embedding-ada-002 (Neelakantan et al., 2022) for Amazon-books and Yelp, and BERT (Devlin, 2018) for Google-reviews. All experiments are conducted on NVIDIA A40/A100 GPUs with 40GB memory.
Figure 3: Ablation study on model variants.
5.2 Performance Comparison
Overall Comparison. The performance comparison across all models is presented in Table 1. Our HERec consistently outperforms the baselines in both utility and diversity metrics, achieving up to $5.49\%$ improvement in utility metrics and $11.39\%$ increase in diversity metrics. Notably, most baselines struggle to achieve strong performance in both utility and diversity simultaneously. For instance, while HICF performs second-best in utility metrics on the Amazon-books dataset, it lacks the same advantage in diversity metrics, where several baselines achieve better results. This highlights the superiority of our model, which achieves state-of-the-art performance in both utility and diversity: a feat not previously attained by any baseline.
Baselines with Feature Enhancement. To ensure a fair comparison, we further enhance Euclidean baselines by integrating feature information. Specifically, we align semantic embeddings with the original Euclidean embeddings following the structure in Section 4.2, but in Euclidean space. The results are presented in Table 2, with enhanced baselines denoted by “+”. The data shows that our HERec still surpasses all baselines in both utility and diversity. Interestingly, while the feature-enhanced versions of the Euclidean baselines generally improve in utility metrics, they do not exhibit a stable improvement in diversity metrics. This aligns with our theoretical analysis in Section 4.4, which demonstrates the advantages of our alignment in the hyperbolic space.
Additionally, to validate the effectiveness of our alignment framework in capturing semantics, we designed a profile decoder module to reconstruct user and item profiles from the learned hyperbolic embeddings. The results demonstrate that profiles reconstructed from our model’s embeddings exhibit higher alignment with the originals, confirming the model’s ability to effectively capture underlying semantic meaning. Detailed methodology and experimental results are provided in Appendix B.
Figure 4: Performance analysis across hierarchical structure layers.
Figure 5: Layer-wise norms within the hierarchy structure.
Performance on Head and Tail Items. To underscore the effectiveness of semantic alignment, particularly for cold-start (tail) items, we conducted experiments measuring utility performance on both head and tail items. In our context, head items (H20) refer to the top $20\%$ of items with the highest number of interactions in the training dataset, representing the most popular items. The remaining items are categorized as tail items (T80), which have limited interaction history, representing less frequently engaged items.
5.3 Ablation Study